# Leveraging Label Information for Stealthy Data Stealing in Vertical Federated Learning

### v1

@@ Abstract.
Launching effective malicious attacks in VFL presents unique challenges: 1) Firstly, given the distributed nature of clients’ data features and models, each client rigorously guards its privacy and prohibits direct querying, complicating any attempts to steal data; 2) Existing malicious attacks alter the underlying VFL training task, and are hence easily detected by comparing the received gradients with the ones received in honest training. To overcome these challenges, we develop DMAVFL, a novel attack strategy that evades current detection mechanisms. The key idea is to integrate a discriminator with auxiliary classifier that takes a full advantage of the label information (which was completely ignored in previous attacks): on one hand, label information helps to better characterize embeddings of samples from distinct classes, yielding an improved reconstruction performance; on the other hand, computing malicious gradients with label information better mimics the honest training, making the malicious gradients indistinguishable from the honest ones, and the attack much more stealthy. Our comprehensive experiments *demonstrate that DMAVFL significantly outperforms existing attacks, and successfully circumvents SOTA defenses for malicious attacks. Additional ablation studies and evaluations on other defenses further underscore the robustness and effectiveness of DMAVFL.
ccs: Security and privacy Distributed systems securityccs: Computing methodologies Machine learning
1. Intoduction
@@
Federated learning (FL) is an emerging privacy-preserving collaborative learning paradigm, which allows multiple distributed clients to securely train a model without sharing private data (Yang et al., 2019b). FL can be categorized into horizontal FL (HFL) and vertical FL (VFL) (Liu et al., 2024). In HFL, different clients possess distinct sets of training samples, yet all samples share the same feature space. In contrast, VFL is applicable to scenarios where clients have the same sample identifiers but possess disjoint feature spaces, e.g., an finance company may hold investment records for a set of customers, while a bank may have the expenditure details for the same set of customers. Due to its capability to integrate diverse data sources, VFL is increasingly gaining attention, and has been applied in various fields including finance (Chen et al., 2020), healthcare (Poirot et al., 2019), and recommendation systems (Yang et al., 2020).
@@
In VFL (Liu et al., 2022; Chen et al., 2020; Li et al., 2023; Thapa et al., 2022; Feng and Yu, 2020; Wu et al., 2020; Yang et al., 2019a), an active client holds both labels and partial features, while multiple passive clients hold disjoint features for the same samples. VFL typically employs a split learning framework (Vepakomma et al., 2018; Abuadbba et al., 2020; Thapa et al., 2022; Kim et al., 2020), and utilizes separate models to process these disjoint features. Specifically, a split VFL system comprises one top model and several bottom models: each passive client operates a bottom model, whereas the active client manages both a bottom model and a top model (see Figure 1). During the training phase, for each batch of training samples, all passive and active clients compute embeddings using their respective bottom models; these embeddings are then transmitted to the active client, who outputs the predictions with the top model and computes the loss from the labels. Following this, the gradients for embeddings are sent back to all clients to update their bottom models.
@@
Figure 1. Illustration of a VFL system with one active client and two passive clients. The area enclosed by the red dashed lines contains the information accessible to and the actions taken by the active client. In a data reconstruction attack, the malicious active client intends to recover the private features of the target passive clients.
@@
While VFL emphasizes privacy by sharing embeddings instead of raw data, emerging data reconstruction attacks have revealed potential vulnerabilities that could lead to privacy leakage within VFL models. Such attacks are categorized into honest-but-curious (HBC) and malicious attacks. HBC adversaries stealthily extract private features while adhering to the protocol, whereas malicious adversaries actively manipulate the protocol to steal private features. Notable methods like GRNA (Luo et al., 2021) and GIA (Jiang et al., 2022) focus on HBC scenarios. In GRNA, under the assumption that the adversary can query the passive clients’ models, target features could be inferred by training a generative regression network to learn correlations between target features and the adversary’s partial features, thus enabling generation of target features. GIA (Jiang et al., 2022), presupposing the adversary can inject adversarial features to the target clients, employs optimization techniques to reconstruct the target features. A malicious data reconstruction attack in VFL, called AGN, is developed in (Vu et al., 2023), where the adversary transmits malicious gradients to passive clients to induce more significant feature leakage.
@@
A class of closely related attacks have also been developed for split learning (SL), which can be viewed as a special case of VFL. In SL, there is a clear division of roles where an active client, holding a top model and labels, but without partial features, collaborates with a single passive client who manages a bottom model and all data samples with complete features. In HBC scenario, attacks like PCAT (Gao and Zhang, 2023) and SDAR (Zhu et al., 2023) use the top model to train a shadow model on an auxiliary dataset, which is subsequently used to train a decoder for data reconstruction. In the malicious scenario, FSHA (Pasquini et al., 2021) manipulates the training objective, generating malicious gradients to achieve more accurate data reconstruction.
@@
While malicious attacks could be more detrimental to privacy leakage in VFL, research works on malicious attacks remains limited. We identify two primary challenges to launch effective malicious attacks: 1) Distributed features and limited model access. The heterogeneous nature of feature sources across distributed clients, together with the lack of access to passive clients’ local models, makes it difficult for the adversary to infer target features using its own partial features; 2) Powerful defenses. Existing malicious attacks, such as FSHA and AGN, replace the honest training task with a malicious task, causing the gradients returned to passive clients to appear different from those in honest training, which renders these attacks easily detectable. Defensive strategies in SL, such as SplitGuard (Erdogan et al., 2022) and Gradient Scrutinizer (Fu et al., 2023), monitor incoming gradients to identify potential attacks. They are readily adapted to the VFL scenario to significantly diminish the success of malicious data reconstruction attacks.
@@
In response to these challenges, we investigate whether a malicious attacker in VFL can circumvent defensive measures to maintain effective data reconstruction. We answer this question affirmatively, by developing a novel defense-penetrating malicious data reconstruction attack on VFL (referred to as DMAVFL). DMAVFL begins with the adversary pretraining an encoder, a decoder, and its bottom model to achieve high reconstruction performance on an auxiliary data. Following this, the adversary freezes the encoder and bottom model, and trains a discriminator with auxiliary classifier (DAC) (Hou et al., 2022) that integrates labels information to generate malicious gradients transmitted to passive clients. These malicious gradients can transfer embeddings distribution from the encoder to passive clients’ models. In data reconstruction phase, the trained decoder is leveraged to reconstruct target features from the target clients, as these passive clients’ models are meticulously guided to mimic the encoder by DAC.
@@
To address challenge 1, DMAVFL captures correlations between the adversary’s partial features and the target features without direct queries or data injections, using DAC to generate malicious gradients that indirectly guide the passive clients’ model to leak feature information. To handle challenge 2, we use DAC to incorporate label information into the computation of the malicious loss function, generating malicious gradients almost indistinguishable from the gradients received in honest training, which helps DMAVFL to effectively penetrate current defenses. Additionally, these label information helps to better transfer the embedding distribution from the encoder to the passive clients’ models compared with a traditional discriminator, and provide extra information for achieving a smaller reconstruction error. As a result, over extensive empirical evaluations, DMAVFL is shown to achieve better reconstruction performance than SOTA attacks, and successfully circumvents all existing defenses against malicious attacks. Further ablation studies and evaluations against other defenses underscore the robustness and effectiveness of DMAVFL.
@@
    We introduce DMAVFL, a novel defense-penetrating malicious data reconstruction attack in VFL, designed to breach existing security measures and effectively steal private features;
  2. (2)
    Unlike traditional approaches that utilize a discriminator, DMAVFL incorporates training a DAC, which helps to 1) significantly enhance the effectiveness in embedding distribution transfer and attack performance; and 2) make the malicious gradients indistinguishable from the ones with honest training;
  3. (3)
    We rigorously evaluate the attack performance of DMAVFL over four representative datasets, demonstrating that it circumvents all current defenses against malicious data reconstruction, and outperforms all SOTA attacks on VFL (and the ones adapted from SL), with and without defenses.
2. Background and Related works
@@
In this section, we review SOTA data reconstruction attacks on VFL, and existing defense mechanisms for these attacks.
2.1. Data reconstruction attacks on VFL
@@
Although VFL safeguards feature privacy by exchanging embeddings and intermediate gradients instead of raw features, recent research has revealed its vulnerability to data reconstruction attacks that can lead to private feature leakage. In HBC setting, the attacks described in (Luo et al., 2021) and (Yang et al., 2023) assume that an HBC adversary can query the passive clients’ bottom model. During the inference phase, the adversary covertly trains a generative model that is fed random vectors and partial data features available to the adversary, thereby producing synthetic inputs. By minimizing the discrepancy of the bottom models’ output between synthetic inputs and target features, the generative model is trained to approximate these features accurately. Notably, while the adversary in (Luo et al., 2021) can access the gradients of the target bottom model directly, the adversary in (Yang et al., 2023) employs zeroth-order optimization to estimate these gradients, thereby bypassing the need for direct gradient access.
@@
The approach in (Jiang et al., 2022), GIA, operates under a black-box scenario where the adversarial active client lacks direct access to the passive clients’ bottom model. Despite this limitation, the adversary can strategically insert a batch of auxiliary data features into the datasets of the passive clients and subsequently collect the resulting output embeddings. The adversary then trains a shadow model using the collected auxiliary features and embeddings pairs. After training, the adversary optimizes noise to closely approximate the target features by querying this shadow model. It is essential to emphasize that these attack methodologies, while effective in reconstructing private data, depend on the adversary having either direct access to the passive clients’ models or the capability to inject data into these clients.
@@
In malicious setting, the attack method, AGN (Vu et al., 2023), utilizes a generator to create fake target features from passive clients’ embeddings. These fake features are then evaluated by a discriminator, which distinguishes between real and fake features to generate malicious gradients. These gradients are sent back to the passive clients, aiming to train embeddings that facilitate the reconstruction of target features. Despite the innovative approach of AGN, our evaluations in the section of experiments reveal that the data reconstructed by this method is relatively coarse and remains susceptible to existing defense mechanisms.
@@
Data reconstruction attacks on split learning. A parallel strand of privacy attacks on SL is adaptable to VFL with certain modifications. Unlike VFL which involves multiple clients, SL typically includes only one client and a server managing the bottom and top models, respectively, which can be viewed as a special case of VFL. In SL, adversarial servers are categorized as either honest-but-curious (HBC) or malicious.
@@
For HBC settings, He et al. pioneer privacy attacks on SL in (He et al., 2019), exploring various levels of adversarial knowledge. In a black-box scenario, where the adversary lacks the knowledge of the model but can make queries, an inverse model is used to reconstruct private features. In scenarios without query capabilities, the adversary utilizes a shadow model trained with auxiliary data under the guidance of the top model. Similar with GIA, the shadow model is then used to reconstruct target features through optimization. Unsplit (Erdoğan et al., 2022) presents a scenario where the adversary knows the bottom model’s structure but lacks auxiliary data. This method involves optimizing synthetic features and the shadow model concurrently to reduce disparities between real and synthetic features’ embeddings. PCAT (Gao and Zhang, 2023) depicts an adversary with access to a batch of auxiliary data but unaware of the bottom model’s structure and parameters. The strategy involves covertly training a shadow model with the auxiliary data, guided by the top model in each training round. Then an inverse model is trained on the shadow model for data reconstruction. SDAR (Zhu et al., 2023), assuming knowledge of the bottom model’s architecture, utilizes auxiliary data and adversarial regularization to train both a shadow model and its inverse model. SDAR employs two discriminators for adversarial regularization: one differentiates between the embeddings of synthetic and genuine features, while the other identifies the data’s origin, whether it comes from the inverse model or the auxiliary data. For malicious setting, the attack method, FSHA (Pasquini et al., 2021), replaces the top model with a discriminator for aligning the bottom model with a local encoder. Subsequently, an decoder trained on the encoder can directly apply to reconstruct target features.
2.2. Defenses against data reconstruction attacks
@@
As data leakage risks in VFL escalate, a variety of defensive strategies have been developed to defend data reconstruction attacks. These defenses, particularly against an HBC adversary, can be broadly classified into cryptographic methods, perturbation-based methods, and adversarial training methods. Cryptographic strategies include homomorphic encryption (HE)(Xu et al., 2021; Gong et al., 2023; Fu et al., 2022), secret sharing (SS)(Li et al., 2023; Huang et al., 2023), differential privacy (DP)(Wang et al., 2020; Chen et al., 2020), and hashing(Qiu et al., 2022). Employed to protect the privacy of individual embeddings in VFL, these methods, while effective, often lead to significant computation and communication overheads. Perturbation methods involve adding noise to embeddings to disturb potential reconstruction attacks, as detailed in (He et al., 2020). Another variation treats noise as trainable parameters, aiming at minimizing privacy leakage without significant degradation of model performance, as explored in (Mireshghallah et al., 2021). While these methods can mitigate data reconstruction attacks, they might reduce model accuracy. Adversarial training techniques, as discussed in (Sun et al., 2021; Vepakomma et al., 2020), train models in such a way that embeddings are less likely to leak raw features information. Although these methods are effective at reducing features leakage, they increase the computational burden and may affect the pefromance of the intended learning tasks.
@@
As malicious attacks such as FSHA and AGN substitute the intended training task, causing significant private features leakage, recent advancements in defense methods, notably SplitGuard (Erdogan et al., 2022) and Gradient Scrutinizer (Fu et al., 2023), have been developed to specifically address these types of attacks. SplitGuard operates by requiring passive clients to intermittently replace labels in a batch with randomized labels, referred to as fake batches, during training. The idea behind SplitGuard is that if the top model is genuinely engaged in learning the intended task, the angles and distances between the gradients from fake and regular batches will differ significantly from those between two regular batches. Consequently, passive clients can detect anomalies by analyzing these differences in gradients. Unlike SplitGuard, Gradient Scrutinizer does not necessitate changes to labels. It operates by collecting and analyzing received batch gradients with identical and differing labels during each training round. The key finding is that during honest training, the gradient distances for embeddings with the different label should be greater than those for embeddings with same labels. By comparing these gradient distances, Gradient Scrutinizer can effectively detect deviations from honest training patterns, indicating potential attacks.
3. Problem Description
3.1. VFL protocol

### v2

# URVFL: Undetectable Data Reconstruction Attack on Vertical Federated Learning

@@ Abstract
Launching effective malicious attacks in VFL presents unique challenges: 1) Firstly, given the distributed nature of clients’ data features and models, each client rigorously guards its privacy and prohibits direct querying, complicating any attempts to steal data; 2) Existing malicious attacks alter the underlying VFL training task, and are hence easily detected by comparing the received gradients with the ones received in honest training. To overcome these challenges, we develop URVFL, a novel attack strategy that evades current detection mechanisms. The key idea is to integrate a discriminator with auxiliary classifier that takes a full advantage of the label information and generates malicious gradients to the victim clients: on one hand, label information helps to better characterize embeddings of samples from distinct classes, yielding an improved reconstruction performance; on the other hand, computing malicious gradients with label information better mimics the honest training, making the malicious gradients indistinguishable from the honest ones, and the attack much more stealthy. Our comprehensive experiments demonstrate that URVFL significantly outperforms existing attacks, and successfully circumvents SOTA detection methods for malicious attacks. Additional ablation studies and evaluations on defenses further underscore the robustness and effectiveness of URVFL. Our code will be available at https://github.com/duanyiyao/URVFL.
publicationid: pubid: Network and Distributedx System Security (NDSS) Symposium 2025 24 - 28 February 2025, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2025.24046 www.ndss-symposium.org
I Intoduction
@@
Federated learning (FL) is an emerging privacy-preserving collaborative learning paradigm, which allows multiple distributed clients to securely train a model without sharing private data [46]. FL can be categorized into horizontal FL (HFL) and vertical FL (VFL) [24]. In HFL, different clients possess distinct sets of training samples, yet all samples share the same feature space. In contrast, VFL is applicable to scenarios where clients have the same sample identifiers but possess disjoint feature spaces, e.g., an finance company may hold investment records for a set of customers, while a bank may have the expenditure details for the same set of customers. Due to its capability to integrate diverse data sources, VFL is increasingly gaining attention and has been applied in various fields, including finance [3], healthcare [29], and recommendation systems [45].
@@
In VFL [23, 3, 22, 35, 8, 41, 44], an active client holds both labels and partial features, while multiple passive clients hold disjoint features for the same samples. VFL typically employs a split learning framework [36, 2, 35, 19], and utilizes separate models to process these disjoint features. Specifically, a split VFL system comprises one top model and several bottom models: each passive client operates a bottom model, whereas the active client manages both a bottom model and a top model (see Figure 1). During the training phase, for each batch of training samples, all passive and active clients compute embeddings using their respective bottom models; these embeddings are then transmitted to the active client, who outputs the predictions with the top model and computes the loss from the labels. Following this, the gradients for embeddings are sent back to all clients to update their bottom models.
@@
Figure 1: Illustration of a VFL system with one active client and two passive clients. The area enclosed by the red dashed lines contains the information accessible to and the actions taken by the active client. In a data reconstruction attack, the malicious active client intends to recover the private features of the target passive clients.
@@
Data reconstruction attacks on VFL: While VFL emphasizes privacy by sharing embeddings instead of raw data, emerging data reconstruction attacks have revealed potential vulnerabilities that could lead to privacy leakage within VFL models. Such attacks are categorized into honest-but-curious (HBC) and malicious attacks. HBC adversaries stealthily extract private features while adhering to the training protocol, whereas malicious adversaries actively manipulate or violate the protocol to steal private features. Notable methods like GRNA [25] and GIA [18] focus on HBC scenarios. In GRNA, under the assumption that the adversary can query the passive clients’ models, target features could be inferred by training a generative regression network to learn correlations between target features and the adversary’s partial features, thus enabling generation of target features. GIA [18], presupposing the adversary can inject adversarial features to the target clients, employs optimization techniques to reconstruct the target features. A malicious data reconstruction attack in VFL, called AGN, is developed in [38], where the adversary transmits malicious gradients to passive clients to induce more significant feature leakage.
@@
A class of closely related attacks have also been developed for split learning (SL), which can be viewed as a special case of VFL. In SL, there is a clear division of roles where an active client, holding a top model and labels, but without partial features, collaborates with a single passive client who manages a bottom model and all data samples with complete features. In HBC scenario, attacks like PCAT [12] and SDAR [50] use the top model to train a shadow model on an auxiliary dataset, which is subsequently used to train a decoder for data reconstruction. In the malicious scenario, FSHA [27] manipulates the training objective, generating malicious gradients to achieve more accurate data reconstruction.
@@
Challenges of existing data reconstruction attacks on VFL: While malicious attacks could be more detrimental to privacy leakage in VFL, research works on malicious attacks remains limited. We identify two primary challenges to launch effective malicious attacks: 1) Distributed features and limited model access. The heterogeneous nature of feature sources across distributed clients, together with the lack of access to passive clients’ local models, makes it difficult for the adversary to infer target features using its own partial features; 2) Powerful detections. Existing malicious attacks, such as FSHA and AGN, replace the honest training task with a malicious task, causing the gradients returned to passive clients to appear different from those in honest training, which renders these attacks easily detectable. SOTA detection strategies for malicious attacks, e.g., SplitGuard [6] and Gradient Scrutinizer [10], monitor incoming gradients to identify potential attacks and significantly diminish the success of malicious data reconstruction attacks.
@@
Our attack URVFL: In response to these challenges, we investigate whether a malicious attacker in VFL can circumvent detection measures to maintain effective data reconstruction. We answer this question affirmatively, by developing a novel undetectable data reconstruction attack on VFL (referred to as URVFL). URVFL begins with the adversary pretraining an encoder, a decoder, and its bottom model to achieve high reconstruction performance on an auxiliary data. Following this, the adversary freezes the encoder and bottom model, and trains a discriminator with auxiliary classifier (DAC) [16] that integrates labels information to generate malicious gradients transmitted to passive clients. These malicious gradients can transfer embeddings distribution from the encoder to passive clients’ models. In data reconstruction phase, the trained decoder is leveraged to reconstruct target features from the target clients, as these passive clients’ models are meticulously guided to mimic the encoder by DAC.
@@
To address challenge 1, URVFL captures correlations between the adversary’s partial features and the target features without direct queries or data injections, using DAC to generate malicious gradients that indirectly guide the passive clients’ model to leak feature information. To handle challenge 2, we use DAC to incorporate label information into the computation of the malicious loss function, generating malicious gradients almost indistinguishable from the gradients received in honest training, which helps URVFL to effectively circumvent current detection strategies. Additionally, these label information helps to better transfer the embedding distribution from the encoder to the passive clients’ models compared with a traditional discriminator, and provide extra information for achieving a smaller reconstruction error. As a result, over extensive empirical evaluations, URVFL is shown to achieve better reconstruction performance than SOTA attacks, and successfully circumvents all existing detections against malicious attacks. Further ablation studies and evaluations against defense strategies underscore the robustness and effectiveness of URVFL.
@@
    We introduce URVFL, a novel undetectable data reconstruction attack on VFL, designed to bypass existing detection measures against malicious attacks and effectively steal private features.
  2. 2.
    We develop a DAC as part of URVFL, which:
    • Significantly enhances the effectiveness of embedding distribution transfer and overall attack performance.
    • Generates malicious gradients that are indistinguishable from those produced during honest training.
  3. 3.
    We conduct rigorous evaluations of URVFL using five representative datasets, demonstrating that it:
    • Circumvents SOTA detection strategies against malicious data reconstruction.
    • Outperforms SOTA data reconstruction attacks on VFL, both with and without detection mechanisms in place.
    • Achieves high-quality reconstruction despite defense methods designed for both malicious and HBC attacks in VFL.
II Background and Related works
@@
In this section, we review SOTA data reconstruction attacks on VFL, and existing detection and defense mechanisms for these attacks.
II-A Data reconstruction attacks on VFL
@@
Although VFL safeguards feature privacy by exchanging embeddings and intermediate gradients instead of raw features, recent research has revealed its vulnerability to data reconstruction attacks that can lead to private feature leakage. In HBC setting, the attacks described in [25] and [47] assume that an HBC adversary can query the passive clients’ bottom model. During the inference phase, the adversary covertly trains a generative model that is fed random vectors and partial data features available to the adversary, thereby producing synthetic inputs. By minimizing the discrepancy of the bottom models’ output between synthetic inputs and target features, the generative model is trained to approximate these features accurately. While the adversary in [25] can access the gradients of the target bottom model directly, the adversary in [47] employs zeroth-order optimization to estimate these gradients, thereby bypassing the need for direct gradient access.
@@
The approach in [18], GIA, operates under a black-box scenario where the adversarial active client lacks direct access to the passive clients’ bottom model. Despite this limitation, the adversary can strategically insert a batch of auxiliary data features into the datasets of the passive clients and subsequently collect the resulting output embeddings. The adversary then trains a shadow model using the collected auxiliary features and embeddings pairs. After training, the adversary optimizes noise to closely approximate the target features by querying this shadow model. It is essential to emphasize that these attack methodologies, while effective in reconstructing private data, depend on the adversary having either direct access to the passive clients’ models or the capability to inject data into these clients.
@@
In malicious setting, the attack method, AGN [38], utilizes a generator to create fake target features from passive clients’ embeddings. These fake features are then evaluated by a discriminator, which distinguishes between real and fake features to generate malicious gradients. These gradients are sent back to the passive clients, aiming to train embeddings that facilitate the reconstruction of target features. Despite the innovative approach of AGN, our evaluations in the section of experiments reveal that the data reconstructed by this method is relatively coarse and remains susceptible to existing defense mechanisms.
@@
Data reconstruction attacks on split learning. A parallel strand of privacy attacks on SL is adaptable to VFL with certain modifications. Unlike VFL which involves multiple clients, SL typically includes only one passive client and an active client managing the bottom and top models, respectively, which can be viewed as a special case of VFL.
@@
For HBC settings, He et al. pioneer privacy attacks on SL in [14], exploring various levels of adversarial knowledge. In a black-box scenario, where the adversary lacks the knowledge of the bottom model but can make queries, an inverse model is used to reconstruct private features. In scenarios without query capabilities, the adversary utilizes a shadow model trained with auxiliary data under the guidance of the top model. Similar with GIA, the shadow model is then used to reconstruct target features through optimization. Unsplit [7] presents a scenario where the adversary knows the bottom model’s structure but lacks auxiliary data. This method involves optimizing synthetic features and the shadow model concurrently to reduce disparities between real and synthetic features’ embeddings. PCAT [12] depicts an adversary with access to a batch of auxiliary data but unaware of the bottom model’s structure and parameters. The strategy involves covertly training a shadow model with the auxiliary data, guided by the top model in each training round. Then an inverse model is trained on the shadow model for data reconstruction. SDAR [50], assuming knowledge of the bottom model’s architecture, utilizes auxiliary data and adversarial regularization to train both a shadow model and its inverse model. SDAR employs two discriminators for adversarial regularization: one differentiates between the embeddings of synthetic and genuine features, while the other identifies the data’s origin, whether it comes from the inverse model or the auxiliary data. For malicious setting, the attack method, FSHA [27], replaces the top model with a discriminator for aligning the bottom model with a local encoder. Concurrently, an decoder trained on the encoder can apply to reconstruct target features.
II-B Defenses and detections against data reconstruction attacks
@@
Defenses: As data leakage risks in VFL escalate, a variety of defensive strategies have been developed to defend data reconstruction attacks. These defenses, particularly against both HBC and malicious adversary, can be broadly classified into cryptographic methods, perturbation-based methods, and adversarial training methods. Cryptographic strategies include homomorphic encryption (HE)[43, 13, 9], secret sharing (SS)[22, 17], and hashing[30]. Employed to protect the privacy of individual embeddings in VFL, these methods, while effective, often lead to significant computation and communication overheads. Perturbation methods involve adding noise to embeddings to disturb potential reconstruction attacks, as detailed in [15, 39, 3]. Another variation treats noise as trainable parameters, aiming at minimizing privacy leakage without significant degradation of model performance, as explored in [26]. While these methods can mitigate data reconstruction attacks, they might reduce model accuracy significantly. Adversarial training techniques, as discussed in [33, 37], train models in such a way that embeddings are less likely to leak raw features information. Although these methods are effective at reducing features leakage, they increase the computational burden and may affect the pefromance of the intended learning tasks.
@@
Detections against malicious attacks: As malicious attacks such as FSHA and AGN substitute the intended training task, causing significant private features leakage, recent advancements in detection methods, notably SplitGuard [6] and Gradient Scrutinizer [10], have been developed to specifically address these types of attacks. SplitGuard operates by requiring passive clients to intermittently replace labels in a batch with randomized labels, referred to as fake batches, during training. The idea behind SplitGuard is that if the top model is genuinely engaged in learning the intended task, the angles and distances between the gradients from fake and regular batches will differ significantly from those between two regular batches. Consequently, passive clients can detect anomalies by analyzing these differences in gradients. Unlike SplitGuard, Gradient Scrutinizer does not necessitate changes to labels. It operates by collecting and analyzing received batch gradients with identical and differing labels during each training round. The key finding is that during honest training, the gradient distances for embeddings with the different label should be greater than those for embeddings with same labels. By comparing these gradient distances, Gradient Scrutinizer can effectively detect deviations from honest training patterns, indicating potential attacks.
Differences between defenses and detections: While defenses are designed to significantly mitigate the privacy leakage by manipulating the embeddings, i.e., the embeddings are modified to preserve its representation ability as well as the privacy, the detection methods will not change the learning process of embeddings and only observe the variations of exchanged information to detect an attack. Specifically, current detections for data reconstruction attack analyze the difference of gradients between the honest and malicious training to detect the attack. Once there is an attack, the victims can stop the training process to prevent further privacy leakage.
III Problem Description
III-A VFL protocol